OpenStudy (anonymous):

A vertical parabola has vertex (4,1) and Y intercept 5. What are its X intercepts?

OpenStudy (anonymous):

y= A(x-4)^2 + 1 is the general form of such a parabola

OpenStudy (anonymous):

we know that when x=0, y= 5

OpenStudy (anonymous):

so we can find A

OpenStudy (anonymous):

5 = 16A +1 A= (1/4)

OpenStudy (anonymous):

so y= (1/4) (x-4)^2 + 1 x intercepts, set y=0

OpenStudy (anonymous):

(x-4)^2 = -4 , this has no solution! , no x intercepts

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

how did you get -4?